OpenStudy (anonymous):

2 log x - log (x+42) =0

OpenStudy (anonymous):

log x + log x - log(x-42) = 0 log (x²/x-42) = 0 x²/x-42 = e⁰

OpenStudy (anonymous):

can you guess the next steps?

OpenStudy (anonymous):

Dont you mean x^2/x-42= 10^0?

OpenStudy (anonymous):

because log is not base of e

OpenStudy (anonymous):

is it log x to the base e or 10? coz if there´s no number as the base of the log, it means the base is e

OpenStudy (anonymous):

k fine, that would mean x²/x-42 = 10⁰

OpenStudy (anonymous):

Yeah I can guess the next steps. Thank you

OpenStudy (anonymous):

yours truly would appreciate you deeply if you could part with a medal

OpenStudy (anonymous):

makes you money?

OpenStudy (anonymous):

TYSM!!